# TermSight: Making Service Contracts Approachable

### v1

@@ F.4. Semi-structured Interview Questions
How can ServiceX use photographs you post for advertisements or promotions? ServiceX can use photos you post in ads without your permission and is not obligated to attribute you as the creator. [Social Media] “…you grant us a worldwide, royalty-free, perpetual, irrevocable, non-exclusive, transferable, and sublicensable license to use, copy, modify, adapt, prepare derivative works of, distribute, store, perform, and display Your Content and any name, username, voice, or likeness provided in connection with Your Content in all media formats and channels now known or later developed anywhere in the world…you irrevocably waive any claims and assertions of moral rights or attribution with respect to Your Content.” (User Agreement)
ServiceX licenses the content you post to third parties. Which of the following is a third-party licensee of user content on ServiceX allowed to do? Train AI models with user content without user consent. [Social Media] “For example, this license includes the right to use Your Content to train AI and machine learning models” (User Agreement) Our licensees cannot: Combine third-party data with ServiceX public content to target ads; Perform background checks; Track, alert, monitor, or investigate sensitive events (for example, protests or rallies) or sensitive organizations (for example, unions or activist groups)…” (Public Content Policy)
You bought a pair of shoes from ServiceY. Before accepting the delivery, you tried the shoes on but found the shoes small for your foot. What can you do to return the item and guarantee a refund? ServiceY does not accept “item doesn’t fit” as a valid reason for return. [E-commerce] “Trades, offline transactions, items that do not fit, or orders where the Buyer changed their mind are not covered by ServiceY Protect. These types of purchases are considered final sale and not eligible for return.” (Return Policy)
According to ServiceY’s authentication policy, when you buy an item over 500\$, ServiceY will first authenticate the item before shipping it to you. If ServiceY decides that the item is fake, what would happen? The item will be disposed. You will be refunded. [E-commerce] “If we reasonably believe an item to be counterfeit, we will provide a refund for the full cost of the item and any accompanying shipping costs to the buyer … if an item in question is deemed counterfeit, we shall ensure disposal of the item at our discretion.” (Counterfeit Items Policy)
Table 1. Example multiple choice comprehension questions: The questions were designed to be difficult and can’t be directly answered by only reading the Summary Snippets. These questions can also require knowledge about information from multiple positions of the same policy or information across different policies.
@@
Figure 19. The figure shows a forest plot of coefficients in the model, corresponding to treatment and control, for each of the experience questions. Each line of the plot shows a 94% High Density Interval (HPDI) for the coefficient. The inner, thicker line represents the 50% HPDI. The results show a significant effect of the TermSight interface on the user experience for every question. Since the 94% HPDI for each pair (control, treatment) for every question do not overlap with each other, we should expect a significant effect of the treatment on the user experience. E1–E7 refer to the 7 experience questions specified in Appendix F.3.
@@
Figure 20. The figure shows a forest plot of coefficients in the model, corresponding to treatment and control, for each of the experience questions. The $x$-axis is on the logistic scale, with +3 corresponding to a 0.95 probability value on the outcome scale (-3 corresponds to 0.05). Each line of the plot shows a 94% High Density Interval (HPDI) for the coefficient. The inner, thicker line represents the 50% HPDI. The 94% HPDI intervals for each treatment–control pair, when the question and service type are fixed, show significant overlap. This suggests no significant effect of the treatment on the user’s comprehension. Note that the comprehension questions Q1–Q6 are different for the two service types and are included in the supplemental materials.
@@
Prompt for Obtaining Summary Snippets and Information Snippets Summarize the input section of the terms of service into concise bullet points (less than 12 words) in plain language. When adjacent paragraphs or sentences share a similar or related theme, only output 1 single bullet point. For each bullet point summary, include the full-text reference to the original passage in {} and don’t use "..." to reduce text in the reference. When outputting a reference, don’t change anything in the original text such as spaces and newlines. There can be multiple sentence or paragraph references to a single summary. The references to summary should cover the original text
Example output format: {EXAMPLE OUTPUT}
Input: {INPUT TEXT CHUNK}
Figure 21. Prompt for obtaining Summary Snippets and Information Snippets
@@
Prompt for Classifying Power Classify the input term from a Terms of Service agreement based on the power relationship and benefit between the service and the user. Use the following categories (Bad, Neutral, Good):
@@
- Blocking first party cookies may limit your ability to use the service
@@
Output format in JSON: {"Category": "Service/Neutral/User", "Explanation": "explanation of output" }
@@
Figure 22. Prompt for classifying power balance of the Information Snippets
Prompt for Classifying Relevance For the input term from a Terms of Service, assign a relevance rating (High/Low) of the input term with respect to the user persona. Output "Low" for low-relevance terms.
Relevance rating:
@@
Output format in JSON: {"Relevance": "Low/High", "Explanation": "explanation of output" }
@@
Figure 23. Prompt for classifying the relevance of Information Snippets to the user persona
Prompt for Identifying Unfamiliar Phrases You are a helpful assistant who extracts words or multi-word phrases in the input section of Terms of Service that a high schooler might not know the meaning of. Jargon refers to domain-specific terminologies that a lay user might not know about.
@@
Output format in JSON: {"Jargon": []}
@@
Prompt for Identifying Vague Phrases You are a helpful legal assistant who extracts vague terms (can have multiple words in one term) in the input section of Terms of Service. A vague term refers to information that is vaguely abstracted without a clear definition provided in the section.
@@
Output format in JSON: {"Vague": []}
@@
Figure 25. Prompt for identifying vague phrases in a text chunk.
Prompt for Generating Phrase Definitions Use information in the retrieved context to provide a definition of the user-selected phrase or term. Avoid using long sentences. For example, if the user-selected term is "information", define what the term "information" includes and refers to, such as: location data, interaction data, profile data, etc. The output definition should be specific and straight to the point, don’t include language that doesn’t contribute to the definition such as ’in the given context’. Output the string list of reference ids (["ref1", ...]) used to generate the definition under "References". If the definition of the phrase is not specified in the retrieved context, output a definition of what the phrase might mean and output an empty array for "References".
@@
Output format in JSON: {"Definition": "", "References": ["ref1", "ref2", "ref3"]}
@@
Question: What does {INPUT PHRASE} refers to?
@@
Figure 26. Prompt for generating in context phrase definitions
Prompt for Generating Scenarios Tell a concise what-if scenario or example in less than 50 words to demonstrate the meaning and potential implications of the user-selected phrase based on the context around the user-selected phrase. The scenario/example should be relevant to the below user persona using {an E-commerce platform of used items / a Social Media platform}.
@@
Output format in JSON: {"Story": ""}
@@
Figure 27. Prompt for generating scenarios to contextualize the meaning and potential implications of a phrase
Prompt for Generating Answers to User Questions You are an assistant for question-answering tasks. Use information in the retrieved context to answer the user’s question in less than 5 sentences. Output the string list of reference ids (["ref1", ...]) used to generate the definition under "References". If the definition of the phrase is not specified in the retrieved context, output a definition of what the phrase might mean and output an empty array for "References".
@@
Output format in JSON: {"Answer": "", "References": ["ref1", "ref2", "ref3"]}
@@
Figure 28. Prompt for generating answers to user questions

### v2

@@ Appendix G Supplemental Analysis for Comprehension Outcomes
Relevant Clause: “…you grant us a worldwide, royalty-free, perpetual, irrevocable, non-exclusive, transferable, and sublicensable license to use, copy, modify, adapt, prepare derivative works of, distribute, store, perform, and display Your Content and any name, username, voice, or likeness provided in connection with Your Content in all media formats and channels now known or later developed anywhere in the world…you irrevocably waive any claims and assertions of moral rights or attribution with respect to Your Content.” (User Agreement)
Possible Room for Interpretation: Though the original answer best matches the relevant clause. The clause does not explicitly mention advertisement. This may leave room for interpretation on whether the clause applies to ads.
We use the same Bayesian model as §F.3, except that there are 4 comprehension questions for each service that are analyzed (i.e., K=4). Out of the 4 questions for each service, participants on average scored 1.5 ($\sigma=1.0$) when using TermSight and 1.6 ($\sigma=1.1$) when using the baseline interface. Our Bayesian model analysis reveals no significant differences in the comprehension scores across interface conditions. The posterior distribution of the contrast between treatment and control on comprehension is centered around zero and overlaps with ROPE. When the question ID and service type are fixed, there are significant overlaps between the 94% HPDI for each treatment–control pair, suggesting no significant differences for every question. These findings matches with our original findings. The figures for this additional analysis can be found in the supplemental materials (e.g., forest plots).
@@
Figure 14. The figure shows a forest plot of coefficients in the model, corresponding to treatment and control, for each of the experience questions. Each line of the plot shows a 94% High Density Interval (HPDI) for the coefficient. The inner, thicker line represents the 50% HPDI. The results show a significant effect of the TermSight interface on the user experience for every question. Since the 94% HPDI for each pair (control, treatment) for every question does not overlap with each other, we should expect a significant effect of the treatment on the user experience. E1–E7 refer to the 7 experience questions specified in Appendix E.5.
@@
The figure shows a forest plot of coefficients in the model, corresponding to treatment and control, for each of the 7 experience questions. Each line of the plot shows a 94% High Density Interval (HPDI) for the coefficient. The inner, thicker line represents the 50% HPDI. There are a total of 14 lines (7 pairs of 2) in the plot. The results show a significant effect of the TermSight interface on the user experience for every question. Since the 94% HPDI for each pair (control, treatment) for every question does not overlap with each other, we should expect a significant effect of the treatment on the user experience. The 7 experience questions are specified in Appendix E.5.
Figure 15. The figure shows a forest plot of coefficients in the model, corresponding to treatment and control, for each of the comprehension questions. The $x$-axis is on the logistic scale, with +3 corresponding to a 0.95 probability value on the outcome scale (-3 corresponds to 0.05). Each line of the plot shows a 94% High Density Interval (HPDI) for the coefficient. The inner, thicker line represents the 50% HPDI. The 94% HPDI intervals for each treatment–control pair, when the question and service type are fixed, show significant overlap. This suggests no significant effect of the treatment on the user’s comprehension. Note that the comprehension questions Q1–Q6 are different for the two service types and are included in the supplemental materials.
The figure shows a forest plot of coefficients in the model, corresponding to treatment and control, for each of the comprehension questions. There are a total of 24 lines (12 pairs of 2). The x-axis is on the logistic scale, with +3 corresponding to a 0.95 probability value on the outcome scale (-3 corresponds to 0.05). Each line of the plot shows a 94% High Density Interval (HPDI) for the coefficient. The inner, thicker line represents the 50% HPDI. The 94% HPDI intervals for each treatment–control pair, when the question and service type are fixed, show significant overlap. This suggests no significant effect of the treatment on the user’s comprehension for every question. Note that the comprehension questions Q1–Q6 are different for the two service types and are included in the supplemental materials.
Figure 16. The figure shows the forest plots of coefficients in the model for user experience, comprehension, and recall with respect to the ordering. Each line of the plot shows a 94% High Density Interval (HPDI) for the coefficient. The inner, thicker line represents the 50% HPDI. Since the 94% HPDI for each ordering pair (first, second) overlaps, there is no significant effect of ordering on user experience, comprehension, and recall.
The figure shows the forest plots of coefficients in the model for user experience, comprehension, and recall with respect to the ordering. Each line of the plot shows a 94% High Density Interval (HPDI) for the coefficient. The inner, thicker line represents the 50% HPDI. Since the 94% HPDI for each ordering pair (first, second) overlaps, there is no significant effect of ordering for user experience, comprehension, and recall.
Prompt for Obtaining Summary Snippets and Information Snippets
@@
Summarize the input section of the Terms of Service into concise bullet points (less than 12 words) in plain language. When adjacent paragraphs or sentences share a similar or related theme, only output 1 single bullet point. For each bullet point summary, include the full-text reference to the original passage in {} and don’t use ”…” to reduce text in the reference. When outputting a reference, don’t change anything in the original text, such as spaces and newlines. There can be multiple sentences or paragraphs that reference a single summary. The references to summary should cover the original text. Example output format: {EXAMPLE OUTPUT} Input: {INPUT TEXT CHUNK}
Figure 17. Prompt for obtaining Summary Snippets and Information Snippets.
Prompt for obtaining Summary Snippets and Information Snippets. Summarize the input section of the terms of service into concise bullet points (less than 12 words) in plain language. When adjacent paragraphs or sentences share a similar or related theme, only output 1 single bullet point. For each bullet point summary, include the full-text reference to the original passage in and don’t use ”…” to reduce text in the reference. When outputting a reference, don’t change anything in the original text, such as spaces and newlines. There can be multiple sentences or paragraphs that reference a single summary. The references to the summary should cover the original text. Example output format: EXAMPLE OUTPUT Input: INPUT TEXT CHUNK.
Prompt for Classifying Power
@@
Classify the input term from a Terms of Service agreement based on the power relationship and benefit between the service and the user. Use the following categories (Service, Neutral, User):
@@
- Blocking first-party cookies may limit your ability to use the service
@@
Output format in JSON: {”Category”: ”Service/Neutral/User”, ”Explanation”: ”explanation of output” }
@@
Figure 18. Prompt for classifying the power balance of the Information Snippets.
[…]
Prompt for Classifying Relevance
For the input term from a Terms of Service, output a relevance rating (High/Low) of the input term with respect to the user persona.
@@
Output format in JSON: {”Relevance”: ”Low/High”, ”Explanation”: ”explanation of output” }
@@
Figure 19. Prompt for classifying the relevance of Information Snippets to the user persona.
Prompt for classifying the relevance of Information Snippets to the user persona. For the input term from a Terms of Service, assign a relevance rating (High/Low) of the input term with respect to the user persona. Relevance rating: [High]: The term is directly relevant to the user’s usage of the service or what the user cares about. The term applies to the user persona and is necessary for the user to know. [Low]: The term is not relevant to the user’s usage of the service or what the user cares about. The term doesn’t apply to the user persona or is not necessary for the user to know. User Persona: INPUT USER PERSONA Output format in JSON: ”Relevance”: ”Low/High”, ”Explanation”: ”explanation of output” Input: INPUT INFORMATION SNIPPET
Prompt for Identifying Unfamiliar Phrases
You are a helpful assistant who extracts words or multi-word phrases in the input section of Terms of Service that a high schooler might not know the meaning of. Jargon refers to domain-specific terminologies that a lay user might not know about.
@@
Output format in JSON: {”Jargon”: []}
@@
Prompt for identifying potentially unfamiliar phrases for lay users in a text chunk. You are a helpful assistant who extracts words or multi-word phrases in the input section of Terms of Service that a high schooler might not know the meaning of. Jargon refers to domain-specific terminologies that a lay user might not know about. Example jargon: - legal jargon: indemnity, arbitration, liability - copyright licenses: sublicensable licenses, royalty-free licenses - technical privacy terms: cookies, Ad identifiers, Authentication tokens Return an empty array if the section does not contain jargon. The extracted word should exactly match the original input text with the same capitalization and sequence of words. Output format in JSON: ”Jargon”: [] Input: INPUT TEXT CHUNK
Prompt for Identifying Vague Phrases
You are a helpful legal assistant who extracts vague terms (can have multiple words in one term) in the input section of Terms of Service. A vague term refers to information that is vaguely abstracted without a clear definition provided in the section.
@@
Output format in JSON: {”Vague”: []}
@@
Figure 21. Prompt for identifying ambiguous phrases in a text chunk.
Prompt for identifying ambiguous phrases in a text chunk. You are a helpful legal assistant who extracts vague terms (can have multiple words in one term) in the input section of Terms of Service. A vague term refers to information that is vaguely abstracted without a clear definition provided in the section. Example Vague terms: information, other, some, third parties, most, generally, personal data, others, general, many, various, might, services, certain information Return an empty array if the section does not contain vague terms. The extracted word should exactly match the original input text with the same capitalization and sequence of words. Output format in JSON: ”Vague”: [] Input: INPUT TEXT CHUNK
Prompt for Generating Phrase Definitions
Use information in the retrieved context to provide a definition of the user-selected phrase or term. Avoid using long sentences. For example, if the user-selected term is ”information”, define what the term ”information” includes and refers to, such as: location data, interaction data, profile data, etc. The output definition should be specific and straight to the point; don’t include language that doesn’t contribute to the definition, such as ’in the given context’. Output the string list of reference ids ([”ref1”, …]) used to generate the definition under ”References”. If the definition of the phrase is not specified in the retrieved context, output a definition of what the phrase might mean and output an empty array for ”References”.
@@
Output format in JSON: {”Definition”: ””, ”References”: [”ref1”, ”ref2”, ”ref3”]}
@@
Question: What does {INPUT PHRASE} refer to?
@@
Figure 22. Prompt for generating in-context phrase definitions.
Prompt for generating in context phrase definitions. Use information in the retrieved context to provide a definition of the user-selected phrase or term. Avoid using long sentences. For example, if the user-selected term is ”information”, define what the term ”information” includes and refers to, such as: location data, interaction data, profile data, etc. The output definition should be specific and straight to the point; don’t include language that doesn’t contribute to the definition, such as ’in the given context’. Output the string list of reference ids ([”ref1”, …]) used to generate the definition under ”References”. If the definition of the phrase is not specified in the retrieved context, output a definition of what the phrase might mean and output an empty array for ”References”. Examples: EXAMPLES Output format in JSON: ”Definition”: ””, ”References”: [”ref1”, ”ref2”, ”ref3”] Retrieved Context: RETRIEVED CONTEXT Question: What does INPUT PHRASE refer to? Context around the user-selected phrase: PHRASE CONTEXT
Prompt for Generating Scenarios
Tell a concise what-if scenario or example in less than 50 words to demonstrate the meaning and potential implications of the user-selected phrase based on the context around the user-selected phrase. The scenario/example should be relevant to the below user persona using {an E-commerce platform of used items / a Social Media platform}.
@@
Output format in JSON: {”Story”: ””}
@@
Figure 23. Prompt for generating scenarios to contextualize the meaning and potential implications of a phrase.
Prompt for generating scenarios to contextualize the meaning and potential implications of a phrase. Tell a concise what-if scenario or example in less than 50 words to demonstrate the meaning and potential implications of the user-selected phrase based on the context around the user-selected phrase. The scenario/example should be relevant to the below user persona using an E-commerce platform of used items / a Social Media platform. User Persona: INPUT USER PERSONA Output format in JSON: ”Story”: ”” User selected phrase: INPUT PHRASE Context around the user-selected phrase: PHRASE CONTEXT Definition of user selected phrase: GENERATED DEFINITION
Prompt for Generating Answers to User Questions.
You are an assistant for question-answering tasks. Use information in the retrieved context to answer the user’s question in less than 5 sentences. Output the string list of reference ids ([”ref1”, …]) used to generate the definition under ”References”. If the definition of the phrase is not specified in the retrieved context, output a definition of what the phrase might mean and output an empty array for ”References”.
@@
Output format in JSON: {”Answer”: ””, ”References”: [”ref1”, ”ref2”, ”ref3”]}
@@
Figure 24. Prompt for generating answers to user questions.
Prompt for generating answers to user questions. You are an assistant for question-answering tasks. Use information in the retrieved context to answer the user’s question in less than 5 sentences. Output the string list of reference ids ([”ref1”, …]) used to generate the definition under ”References”. If the definition of the phrase is not specified in the retrieved context, output a definition of what the phrase might mean and output an empty array for ”References”. Examples: EXAMPLES Output format in JSON: ”Answer”: ””, ”References”: [”ref1”, ”ref2”, ”ref3”] Retrieved Context: RETRIEVED CONTEXT Question: USER QUESTION User selected phrase: INPUT PHRASE Context around the user-selected phrase: PHRASE CONTEXT
Four examples of the Phrase Scope’s generated definitions and scenarios for potentially unfamiliar and vague phrases. The four columns describe the type of phrase (Jargon or Vague phrase), the original passage that contains the phrase, the generated definition, and the generated scenario.
Figure 25. The baseline reading interface. Participants can navigate to different policies using the top navigational panel. In place of the Summary Snippets on the left is a table of contents. Participants can click a section header in the table of contents to navigate to the corresponding section.
The baseline system used in the user study. The interface contains three panels: a top navigation panel for switching between different policies, the original text on the right, and the table of contents on the left. Participants could click a section header in the table of contents to navigate to the corresponding section.